OpenStudy (anonymous):

factor the trinomial completely x^2+40x+41 is the answer (the polynomial a prime)?

OpenStudy (anonymous):

Edit: Typo. What are the factors of 41 that can be multiplied to get 40? They are -1 and 41. Therefore, x^2+40x+41= (x + 41)(x - 1) The polynomial is not prime, as its factors are (x + 41) and (x - 1).

OpenStudy (anonymous):

The first sentenced should say added, not multiplied.

OpenStudy (anonymous):

what will make a trinomial a prime?

OpenStudy (anonymous):

If we cannot factor it into elements like (x + some number) or if there is nothing common to all terms in the polynomial, i.e. a number, such as 2, or a variable, such as x.

OpenStudy (anonymous):

@SwiftDrift am i right in this one x^2-13x+40 = (x+5)(x+8)

OpenStudy (anonymous):

Close. You should have minuses instead of pluses. (-5) times (-8) = 40 (-5) plus (-8) = -13

OpenStudy (anonymous):

so it would be (x-5)(x-8)

OpenStudy (anonymous):

Correct, that is the factored for of the polynomial.

OpenStudy (anonymous):

form*

OpenStudy (anonymous):

thanks !

OpenStudy (anonymous):

@SwiftDrift sorry to bother but am i right in this one x^2-2x-15=(x-5)(x+3)

OpenStudy (anonymous):

Very good!

OpenStudy (anonymous):

thanks again :)